# Coherence Length of Electronic Nematicity in Iron-Based Superconductors

### v1

@@ Abstract
Figure 2: LD images in the electronic nematic states of iron-based superconductors. (a) LD image in BaFe2As2 at 55 K measured by linearly polarized light along the orthorhombic axes (crossed arrows). (b) Spatial dependence of normalized order parameter $w$ in BaFe2As2 near a domain boundary. (c) LD image at 45 K and (d) normalized order parameter for FeSe0.9S0.1. The data (open circles) in (b) and (d) are obtained by averaging LD signals in the black square area of (a) and (c) and plotted as a function of the distance along the arrows in (a) and (c), respectively. Red lines represent fitting results to Eq. (2) (red lines). The domain walls are highlighted in green shades.
Figures 2(a) and (c) show the LD images obtained for the cleaved surfaces of Ba122 and FeSe0.9S0.1, respectively. In both cases, clear stripe patterns of LD asymmetry are resolved, which signifies the domain structures of the nematic orbital order below $T_{\rm s}$. The LD signal was taken by the difference in the photoemission intensities of parallel and perpendicular polarized light to the orthorhombic axis. To evaluate the domain wall thickness, we analyze the data in areas where the domain wall interval $S$ is wide enough. We use the Ginzburg-Landau (GL) theory, in which the free energy of the system can be written as [20]
@@
FeSe0.9S0.1 0.29 480 78
[…]
where $\phi(x)$ is the order parameter shown in Fig. 1 and Eq. (2), and $2\sigma\sqrt{2\log 2}$ corresponds to the spatial resolution, which is the full width at half maximum of the Gaussian. In Fig. 2(b), we plot the spatial dependence of the normalized LD order parameter $w=\Phi(x)/D$ near a domain boundary in Ba122. In this material, the LD signal changes sign sharply near the boundary, and the spatial dependence of $w$ can be fitted nicely with Eq. (3). However, we find that the nematic coherence length is resolution limited, namely $\xi_{\mathrm{nem}}\lesssim 92$ nm. In the previous optical measurements of Ba122 systems, the reported typical domain size is around a few hundred nm to a few $\mu$m [17, 28, 29], which results in the condition $S\gg\xi_{\mathrm{nem}}$. Correspondingly, we only observe square-wave-like domain structures rather than a sinusoidal wave.
In contrast, the LD signal in FeSe0.9S0.1 shows a much broader change at the boundary as shown in Fig. 2(d). The fitting gives $\xi_{\mathrm{nem}}=480$ nm, which is much longer than the typical lattice domain wall thickness, as previously found in FeSe and BaFe2(As0.87P0.13)2 [20]. Combining these results with the previous work, we summarize the nematic coherence length $\xi_{\mathrm{nem}}$ for four iron-based superconductors in Table 1. These results show that the coherence length of electronic nematic order varies in these iron-based superconductors. We note that the nematic phase accompanies antiferromagnetism in Ba122 systems but not in FeSe-based materials, implying that the order of spin state is not the key parameter for the coherence length. We also note that the domain wall thickness does not change significantly as a function of temperature below $T_{\rm s}$ [20], although the simple GL free energy discussion predicts $T$-dependent $\xi_{\mathrm{nem}}$. This suggests that additional pinning mechanisms may exist in determining the domain structures.
@@
To gain some insights into this unusual material variation of nematic coherence length, we compare the temperature dependence of resistivity $\rho(T)$ in these iron-based superconductors, as shown in Fig. 3(a). The $\rho(T)$ curves show noticeable anomalies at $T_{\rm s}$, below which the resistivity goes up or down depending on the competition between changes in the density of states and scattering rate. We fit the data above the nematic transition $T_{\rm s}$ to the simple power-law dependence $\rho(T)=\rho_{0}+AT^{n}$, where $\rho_{0}$ is the residual resistivity, $A$ is the coefficient, and $n$ is the temperature exponent. We find that the exponent $n$ varies strongly in these materials, ranging from $n\sim 2$ expected for the standard Fermi-liquid theory of metals to $n\sim 1$, a hallmark of non-Fermi liquid properties. Comparisons between this temperature exponent $n$ and nematic coherence length $\xi_{\mathrm{nem}}$ are indicated in Fig. 3(b), showing a correlation between the non-Fermi liquid properties and long $\xi_{\mathrm{nem}}$ values. The results suggest that the source of deviations from the Fermi-liquid properties above $T_{\rm s}$, which is most likely spin-orbital fluctuations that are relevant to the nematic ordering at $T_{\rm s}$, plays a significant role in enhancing the coherence length of electronic nematic order.
@@
Recent electron-correlation theories based on Aslamazov-Larkin vertex corrections [22, 23, 32] suggest that the electronic nematic order can be driven by the multimode interference effects of spin-orbital fluctuation channels. This highlights the importance of antiferroic fluctuations in stabilizing the nematic orbital order, in which the dynamic susceptibility $\chi(\bm{q})$ is governed by the term $(\bm{q}-\bm{Q})^{-1}$ where $\bm{Q}$ is the nesting vector [22]. It would be then expected that when such fluctuations that drive the nematic order are strong enough to give rise to deviations from the Fermi-liquid behavior, the nematic order may have high stiffness, which results in a long coherence length. Within this framework, the GL coefficients for the nematic order, $a\equiv{\dot{a}}(T-T_{\rm s})$ and $c$ in Eq. (1), have been studied theoretically [23]. Here ${\dot{a}}\ (>0)$ is the temperature slope of $a(T)$ at the transition temperature. The GL theory for Bardeen-Cooper-Schrieffer (BCS) superconductors gives the relations ${\dot{a}}\sim N(0)T_{\rm c}^{-1}$ and $c\sim N(0)(\bm{K}/2)^{-2}$, where $N(0)$ is the density of states at the Fermi level, $T_{\rm c}$ is the critical temperature, and $\bm{K}$ is the reciprocal vector. For the nematic order, in contrast, the relations ${\dot{a}}\ll N(0)T_{\rm s}^{-1}$ and $c\gg N(0)(\bm{K}/2)^{-2}$ can be derived [23]. The obtained small ${\dot{a}}$ explains the wide nematic critical region and the small specific heat jump at $T_{\rm s}$ in iron-based superconductors. The large $c$ coefficient originates from the good nesting between the hole and electron Fermi-surface pockets. Therefore, the nematic domain wall thickness $\xi_{\rm nem}\propto\sqrt{2c/|a|}$ can become much longer than the BCS coherence length $\xi_{\rm BCS}\sim v_{\rm F}/T_{\rm c}$ with the Fermi velocity $v_{\rm F}$. Thus, the nematic coherence length can be strongly enhanced by the underlying antiferroic fluctuations, which are related to the Fermi-surface nesting.

### v2

@@ Abstract
Figure 2: LD images in the electronic nematic states of iron-based superconductors. (a) LD image in BaFe2As2 at 55 K and (b) in FeSe0.9S0.1 at 45 K measured by linearly polarized light along the orthorhombic axes (crossed arrows). (c) Spatial dependence of normalized order parameter $w$ in BaFe2As2 obtained by averaging LD signals in area 1 of (a) near a domain boundary (dashed square). (d) Similar data obtained in area 2 of (a) containing several boundaries (solid square). (e) Spatial dependence of normalized order parameter $w$ in FeSe0.9S0.1 obtained by averaging LD signals in area 1 of (b) near a domain boundary (dashed square). (f) Similar data obtained in area 2 of (b) containing several boundaries (solid square). The data (open circles) in (c)-(f) obtained by averaging LD signals are plotted as a function of the distance along the arrows in (a) and (b). Red lines in (c)-(f) represent fitting results to Eq. (2). The domain walls are highlighted in green shades in (c) and (e).
@@
Figures 2(a) and (b) show the LD images obtained for the cleaved surfaces of Ba122 and FeSe0.9S0.1, respectively. In both cases, clear stripe patterns of LD asymmetry are resolved, which signifies the domain structures of the nematic orbital order below $T_{\rm s}$. The LD signal was taken by the difference in the photoemission intensities of parallel and perpendicular polarized light to the orthorhombic axis. To evaluate the domain wall thickness, we analyze the data in areas where the domain wall interval $S$ is wide enough. We use the Ginzburg-Landau (GL) theory, in which the free energy of the system can be written as [20]
@@
[…]
where $\phi(x)$ is the order parameter shown in Fig. 1 and Eq. (2), and $2\sigma\sqrt{2\log 2}$ corresponds to the spatial resolution, which is the full width at half maximum of the Gaussian. In Fig. 2(c), we plot the spatial dependence of the normalized LD order parameter $w=\Phi(x)/D$ near a domain boundary in Ba122. In this material, the LD signal changes sign sharply near the boundary, and the spatial dependence of $w$ can be fitted nicely with Eq. (3). However, we find that the nematic coherence length is resolution limited, namely $\xi_{\mathrm{nem}}\lesssim 92$ nm. We confirm this resolution-limited changes for several domain walls as shown in Fig. 2(d). In the previous optical measurements of Ba122 systems, the reported typical domain size is around a few hundred nm to a few $\mu$m [17, 28, 29], which results in the condition $S\gg\xi_{\mathrm{nem}}$. Correspondingly, we only observe square-wave-like domain structures rather than a sinusoidal wave.
In contrast, the LD signal in FeSe0.9S0.1 shows a much broader change at the boundary as shown in Figs. 2(e) and (f). Here we fit the data for 11 domain walls simultaneously by varying the interval $S$. As a result, all the data can be fitted successfully by taking the coherence length $\xi_{\mathrm{nem}}=310$ nm as the fitting parameter. This value is not resolution limited, which indicates that the orbital nematic coherence length is much longer than the typical lattice domain wall thickness, as previously found in FeSe and BaFe2(As0.87P0.13)2 [20]. Combining these results with the previous work, we summarize the nematic coherence length $\xi_{\mathrm{nem}}$ for four iron-based superconductors in Table 1. These results show that the coherence length of electronic nematic order varies in these iron-based superconductors. We note that the nematic phase accompanies antiferromagnetism in Ba122 systems but not in FeSe-based materials, implying that the order of spin state is not the key parameter for the coherence length. We also note that the domain wall thickness does not change significantly as a function of temperature except near $T_{\rm s}$, where the analysis becomes difficult [20]. This apparently contradicts the temperature dependent $\xi_{\mathrm{nem}}$ expected in the simple GL free energy discussion based on Eq. (1). However, this may be related to the weakly first-order transition at $T_{\rm s}$ found in some iron-based superconductors [14], which requires additional higher-oder terms of the order parameter in the free energy. In such a case, the divergence of $\xi_{\mathrm{nem}}$ can be avoided. Moreover, the apparent temperature-independent behavior of coherence length can be explained by the strong-coupling effect of the nematic order, which flattens the temperature dependence of energy gap ($\propto 1/\xi_{\mathrm{nem}}$) except near the transition temperature.
@@
FeSe0.9S0.1 0.29 310 78
@@
To gain some insights into the observed unusual material variation of nematic coherence length, we compare the temperature dependence of resistivity $\rho(T)$ in these iron-based superconductors, as shown in Fig. 3(a). The $\rho(T)$ curves show noticeable anomalies at $T_{\rm s}$, below which the resistivity goes up or down depending on the competition between changes in the density of states and scattering rate. We fit the data above the nematic transition $T_{\rm s}$ to the simple power-law dependence $\rho(T)=\rho_{0}+AT^{n}$, where $\rho_{0}$ is the residual resistivity, $A$ is the coefficient, and $n$ is the temperature exponent. We find that the exponent $n$ varies strongly in these materials, ranging from $n\sim 2$ expected for the standard Fermi-liquid theory of metals to $n\sim 1$, a hallmark of non-Fermi liquid properties. Comparisons between this temperature exponent $n$ and nematic coherence length $\xi_{\mathrm{nem}}$ are indicated in Fig. 3(b), showing a striking correlation between the non-Fermi liquid properties and long $\xi_{\mathrm{nem}}$ values. The results suggest that the source of deviations from the Fermi-liquid properties above $T_{\rm s}$, which is most likely spin-orbital fluctuations that are relevant to the nematic ordering at $T_{\rm s}$, plays a significant role in enhancing the coherence length of electronic nematic order.
@@
Recent electron-correlation theories based on Aslamazov-Larkin vertex corrections [22, 23, 32] suggest that the electronic nematic order can be driven by the multimode interference effects of spin-orbital fluctuation channels. This highlights the importance of antiferroic fluctuations in stabilizing the nematic orbital order, in which the dynamic susceptibility $\chi(\bm{q})$ is governed by the term $(\bm{q}-\bm{Q})^{-1}$ where $\bm{Q}$ is the nesting vector [22]. It would be then expected that when such fluctuations that drive the nematic order are strong enough to give rise to deviations from the Fermi-liquid behavior, the nematic order may have high stiffness, which results in a long coherence length. Within this framework, the GL coefficients for the nematic order, $a\equiv{\dot{a}}(T-T_{\rm s})$ and $c$ in Eq. (1), have been studied theoretically [23]. Here ${\dot{a}}\ (>0)$ is the temperature slope of $a(T)$ at the transition temperature. The GL theory for Bardeen-Cooper-Schrieffer (BCS) superconductors gives the relations ${\dot{a}}\sim N(0)T_{\rm c}^{-1}$ and $c\sim N(0)(\bm{K}/2)^{-2}$, where $N(0)$ is the density of states at the Fermi level, $T_{\rm c}$ is the critical temperature, and $\bm{K}$ is the reciprocal vector. For the nematic order, in contrast, the relations ${\dot{a}}\ll N(0)T_{\rm s}^{-1}$ and $c\gg N(0)(\bm{K}/2)^{-2}$ can be derived [23]. The obtained small ${\dot{a}}$ explains the small specific heat jump at $T_{\rm s}$ in iron-based superconductors. The large $c$ coefficient originates from the good nesting between the hole and electron Fermi-surface pockets. Therefore, the nematic domain wall thickness $\xi_{\rm nem}\propto\sqrt{2c/|a|}$ can become much longer than the BCS coherence length $\xi_{\rm BCS}\sim v_{\rm F}/T_{\rm c}$ with the Fermi velocity $v_{\rm F}$. Thus, the nematic coherence length can be strongly enhanced by the underlying antiferroic fluctuations, which are related to the Fermi-surface nesting.